OpenStudy (anonymous):

CHALLENGE QUESTION Solve in positive real number the system of equation : \((2x)^{2013} + (2y)^{2013} + (2z)^{2013} = 3\) and \(xy + yz + zx + 2xyz = 1\)

OpenStudy (anonymous):

Enjoy \(2013\) :)

Parth (parthkohli):

Nothing to do here.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

@ParthKohli nothing to do?

OpenStudy (raden):

x = y = z = 1/2 proving : 1^3 + 1^3 + 1^3 = 3 true 1/4 + 1/4 + 1/4 + 1/4 = 1 true too hohohoho :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

setting up the lagrangian should be enough, wouldn't it?

OpenStudy (anonymous):

I think it would involve "modulus" but I don't know how..

OpenStudy (anonymous):

\[f(x,y,z)=x^{2013}+y^{2013}+z^{2013}-3\times2^{-2013}\\ g(x,y,z)=xy+yz+zx+2xyz\\\;\\ f_x=\lambda g_x\implies2013x^{2012}=\lambda(y+z+2yz)\\ f_y=\lambda g_y\implies2013y^{2012}=\lambda(x+z+2xz)\\ f_z=\lambda g_z\implies2013z^{2012}=\lambda(x+y+2xy)\\ g(x,y,z)=c\implies xy+yz+zx+2xyz=1 \]

OpenStudy (anonymous):

That looks like you have tried something tex work but I would have to go through that first ...

OpenStudy (anonymous):

from 1,3 \[y+z=\frac{2013x^{2012}}{\lambda}-2yz\\ x(y+z)+yz+2xyz=1\implies \frac{2013x^{2013}}{\lambda}-2xyz+yz+2xyz=1\\ \implies\boxed{\lambda yz=2013x^{2013}} \] similarly, others \(\lambda zx,\lambda xy\)

OpenStudy (anonymous):

Gud work @electrokid So, is that your final answer

OpenStudy (anonymous):

thats half the steps.. find out lambda

OpenStudy (anonymous):

Ok so go on.........

OpenStudy (anonymous):

or use the three relations and find out the equations for "x","y" and "z" lambdas cancel out

OpenStudy (anonymous):

what do you men go on? go ahead and plug em up back into the relations..

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

I would after that....... Gud work